Grade 11Physical Chemistry

H2SO4 solution is 500 molal. (density = 1.8 g/mL).Hence its molarity is
(a) 18M
(b) 36M
(c) 9M
(d) 4.50M

RAJIV SHARMA
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
We know that,

molarity=no. of  moles ofsolute/vol of soln. in litre

and

molality=no. of moles of solute/ wt. of solvent.

Dividing both we get the relation as:
M/m = wt. of solvent/vol. of soln.
Further elaborating we get ,
 

molality= Molarity/(density of solution-Molarity*molecular wt of solute)

 

m = M / (ρ - M x mw A) 

Putting the values,

M = m*(ρ - M x mw A) 

  M = 500*(1.8 – M*98)

M = 900 – 49000M

49001M = 900

M = 900/49001 => 0.0183 *1000 => 18.3 approx.